OpenStudy (anonymous):

Give the formula, solve for the indicated variable. F=9/5C+32 for C Please show work

jimthompson5910 (jim_thompson5910):

\[\Large F = \frac{9}{5}C+32\] \[\Large F-32 = \frac{9}{5}C\] \[\Large 5(F-32) = 9C\] \[\Large \frac{5}{9}(F-32) = C\] \[\Large C=\frac{5}{9}(F-32)\]

OpenStudy (anonymous):

jim can you knock out my problem really quick